OpenStudy (lgbasallote):

\[\huge \int \frac{\sqrt{1 + \sqrt x}}x dx\]

OpenStudy (lgbasallote):

it looks tough....

OpenStudy (lgbasallote):

i forgot to add...solve by u-sub

OpenStudy (lgbasallote):

ahh i have an idea \[a = \sqrt x\] \[a^2 = x\] \[2ada = dx\] \[\huge \implies \int \frac{\sqrt{1-a}}{a^2}(2ada)\]

OpenStudy (lgbasallote):

then i suppose i can do this \[\huge \implies 2\int \frac{\sqrt {1-a}}a da\]

OpenStudy (lgbasallote):

then the rest will be easy

OpenStudy (lgbasallote):

thanks